OpenStudy (anonymous):

\[\int_{2}^{6} 2\pi x \sqrt{1-\left(\frac{9}{2}\left(9x-18\right)^{-\frac 12}\right)^2}\] \[2\pi \int_{2}^{6} x \sqrt{1-\frac{18}{4}\left( 9x-18\right)^{-1}}dx\] Where from here? I know...

OpenStudy (anonymous):

x=(y^2/9) +2

OpenStudy (anonymous):

pls edit.

OpenStudy (anonymous):

what do you want me to edit? The original equation was \[y=\sqrt{9x-18}\] or \[x=\frac{y^2}{9} +2\]

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

\[2\pi \int_{2}^{6} x\sqrt{1-\frac{1}{2(x-2)}} dx\]

OpenStudy (anonymous):

rotating about x-axis

OpenStudy (anonymous):

I don't know how to get rid of the \[\sqrt ...\]

OpenStudy (anonymous):